A small personal computer company sells five different computer models at three stores located throughout a large metropolitan area. The inventory of each model on hand in each store is summarized in MATRIX A. MATRIX B summarizes the wholesale and retails values of each model. What is the total wholesale value and retail value of these computers at Store 2? # of Each Type of Computer Model MATRIX A Store 1 Store 2 Store 3 14 12 10 8 MATRIX B Wholesale Retail Model A 200.00 400.00 Model B 210.00 425.00 Model C220.00 450.00 Model D 250.00 500.00 Model E 400.00 $1,000.00 Wholesale Value $6,780, Retail Value $14,250 Wholesale Value $7,730, Retail Value $15,750 Wholesale Value $7,910, Retail Value $16,350 Wholesale Value $8,760, Retail Value $18,600 Wholesale Value $7,440, Retail Value $15,750
Answer #1

1) The Wholesale value at store 2 is calculated as:

(12*200 + 6*210 + 5*220 + 8*250 + 5*400) = (2400+1260+1100+2000+2000) = $ 8,760.

as the wholesale values are given to be so and there are 12,6,5,8,5 numbers of inventories respecively to the models A,B,C,D,E.

2) The Retail value at store 2 is calculated as:

(12*400 + 6*425 + 5*450 + 8*500 + 5*1000) = (4800+2550+2250+4000+5000) = $ 18,600.

as the retail values are given to be so and there are 12,6,5,8,5 numbers of inventories respecively to the models A,B,C,D,E.

So, option 4, Wholesale value = $ 8,760, Retail value =$ 18,600 is the correct one.